How much of an alloy that is 30% copper should be mixed with 400 ounces of an alloy that is 90% copper in order to get an alloy that is 70%?
Answer by macston(5194)

.
T=amount of 30% alloy
.
0.30T+0.90(400)=0.70(T+400)
0.30T+360=0.70T+280
80=0.40T
200=T
.
ANSWER: 200 ounces of 30% alloy should be used.
.